OpenStudy (anonymous):

A chocolate chip cookie recipe asks for two and one half times as much flour as chocolate chips. If three and three quarters cups of flour is used, what quantity of chocolate chips would then be needed, according to the recipe?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@mr570 any idea?

OpenStudy (anonymous):

Ok so. What you're wanting is 2 1/2 times as much flour as chocolate chips for the recipe. You know that you have 3 3/4 cups of flour. You can solve this problem algebraically. What I would do first is change the mixed fractions of 2 1/2 and 3 3/4 into improper fractions. This will give you 2 1/2 = 5/2 and 3 3/4 = 15/4. So once you've done that you can set up an equation (solve algebraically!). The equation will look like this: (amount of flour you have)(a variable [like x or y or whatever]) = (amount of chocolate chips you want). Are you following so far?

OpenStudy (anonymous):

mhm

OpenStudy (anonymous):

Okay. Do you have any questions before I continue?

OpenStudy (anonymous):

not yet.

OpenStudy (anonymous):

continue please

OpenStudy (anonymous):

Okay. So now that you have your equation, let's say you chose x for your variable, you can solve like so: 5/2 x = 15/4. Do you know how to divide by a fraction?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

What you do is instead of dividing by the fraction, you multiply by the fraction's reciprocal. Do you know what a reciprocal is? If not, could you tell me what the parts of a fraction are? Like what is the top half of the fraction called and what is the bottom half of the fraction called?

OpenStudy (anonymous):

Keep in mind I'm just trying to help you learn... Still there?

OpenStudy (anonymous):

Okay just to close this out, look here: